OpenStudy (anonymous):

evaluate the integral using the Fundamental Theorem of Calculus (i.e. antiderivatives) to determine the exact value: integral of (10)/(3x+1) from 0 to 3

OpenStudy (gorv):

\[\int\limits_{0}^{3}\frac{ 10 }{ 3x+1 } *dx\]

OpenStudy (gorv):

\[10*\int\limits_{0}^{3}\frac{ dx }{ 3x+1 }\]

OpenStudy (gorv):

as 10 is constant u can take it out

OpenStudy (gorv):

let 3x+1=u so new limit will be for x=0 u=3*0+1=1 for x=3 u=3*3+1=10 from 1 to 10 now 3x+1=u differntiate 3*dx=du dx=du/3

OpenStudy (gorv):

so now nem integral is \[10*\int\limits_{1}^{10}\frac{ du }{ 3 *u }\ ]take 1/3 out \[\frac{ 10 }{ 3 }* \int\limits_{1}^{10} \frac{ du }{ u }\]

OpenStudy (gorv):

du/u=Ln(u) ln= natural log

OpenStudy (anonymous):

hello @gorv ! so are your writing all of this out it is all coming back to me!! :)

OpenStudy (gorv):

\[\frac{ 10 }{ 3 }\left[ \ln u \right]\] with limi from 1 to 10

OpenStudy (anonymous):

yay i got that as well :)

OpenStudy (gorv):

what ????

OpenStudy (anonymous):

First approaching the problem i couldn't think of how to solve. but as you were explaining, i was able to remember how do to it. but i very much appreciate your help!

OpenStudy (gorv):

ohhhh i thought somthin else lolll